OpenStudy (anonymous):

A car slams on its brakes, coming to a complete stop in 4.0 s. The car was traveling north at 60.0 mph. Calculate the acceleration.

OpenStudy (anonymous):

v=v0+at

OpenStudy (anonymous):

v will be 0 as it end with a stop, v0=60 initial speed

OpenStudy (anonymous):

so substitute and calculate the acceleration, it can be clearly seen that it is in negative direction

OpenStudy (anonymous):

0=60+a4 64 a?

Parth (parthkohli):

Here:\[\rm acceleration = \dfrac{initial ~velocity - final~velocity}{time~gap}\]

OpenStudy (anonymous):

0-60.0/4

Parth (parthkohli):

And how do you simplify that?

OpenStudy (anonymous):

a=-15

OpenStudy (anonymous):

you have just solved

OpenStudy (anonymous):

thanks!